OpenStudy (anonymous):

3/4x=18 how can i divide this equation ?

OpenStudy (anonymous):

multiply by its reciprocal

OpenStudy (jack1):

3/4x = 18 3 = 18 x 4x 3/18 = 4x x = 3/72

OpenStudy (anonymous):